# GnnX-Bench: Unravelling the Utility of Perturbation-based Gnn Explainers through In-depth Benchmarking

### v1

@@ 4.2 Stability
We next examine the stability of the explanations against topological noise, model parameters, and the choice of Gnn architecture. Given a graph , let and be the set of edges in the original and perturbed explanations respectively. We measure stability by computing the Jaccard similarity between the two sets, i.e, . In App. B, we present the impact of the above mentioned factors on other metrics of interest such as sufficiency and explanation size. In addition, we also present impact of feature perturbation and topological adversarial attack in App. B.
@@
Factual-stability against topological noise: Fig. 3 illustrates the Jaccard coefficient as a function of the noise volume. Similar to Fig.2, edge selection for the explanation involves a greedy approach that prioritizes the highest score edges. RCExplainer (executed at ) and PGExplainer consistently exhibit higher stability. This consistent performance reinforces the claim that RCExplainer is the preferred factual explainer. The stability of RCExplainer can be attributed to its strategy of selecting a subset of edges that is resistant to changes, such that the removal of these edges significantly impacts the prediction made by the remaining graph [6]. PGExplainer also incorporates a form of inherent stability within its framework. It builds upon the concept introduced in GNNExplainer through the assumption that the explanatory graph can be modeled as a random Gilbert graph, where the probability distribution of edges is conditionally independent and can be parameterized. This generic assumption holds the potential to enhance the stability of the method. Conversely, TagExplainer exhibits the lower stability than RCExplainer and PGExplainer, likely due to its reliance solely on gradients in a task-agnostic manner [51]. The exclusive reliance on gradients makes it more susceptible to overfitting, resulting in reduced stability.
@@
Factual-Stability against explainer instances: Table 6 presents the stability of explanations provided across three different explainer instances on the same black-box Gnn. A similar trend is observed, with RCExplainer remaining the most robust method, while GnnExplainer exhibits the least stability. For GnnExplainer, the Jaccard coefficient hovers around , indicating significant variance in explaining the same Gnn. This lack of stability may hinder the practical adoption of factual explainers for real-world use cases.
@@
Factual-Stability against Gnn architectures: Finally, we explore the stability of explainers across different Gnn architectures in Table 7, which has not yet been investigated in the existing literature. For each combination of architectures, we assess the stability by computing the Jaccard coefficient between the explained predictions of the indicated Gnn architecture and the default Gcn model. One notable finding is that the stability of explainers exhibits a strong correlation with the dataset used. Specifically, in five out of six datasets, the best performing explainer across all architectures is unique. However, it is important to highlight that the Jaccard coefficients across architectures consistently remain low indicating stability against different architectures is the hardest objective due to the variations in their message aggregating schemes.
@@
| RCExplainer | CF | CLEAR | |||||||
| Dataset / Seeds | |||||||||
| Mutagenicity | 0.96 ±0.06 | 0.96 ±0.04 | 0.98 ±0.03 | 0.90 ±0.09 | 0.89 ±0.1 | 0.89 ±0.11 | OOM | OOM | OOM |
| Proteins | 0.95 ±0.0 | 0.94 ±0.0 | 0.90 +- 0.0 | NA | NA | NA | OOM | OOM | OOM |
| Mutag | 0.98 ±0.03 | 0.98 ±0.03 | 1.0 ±0.0 | 1.0 ±0.0 | 1.0 ±0.0 | 1.0 ±0.0 | 0.55 ±0.01 | 0.53 ±0.01 | 0.54 ±0.02 |
| IMDB-B | 0.99 ±0.01 | 1.0 ±0.0 | 0.99 ±0.01 | 0.96 ±0.05 | 0.95 ±0.06 | 0.94 ±0.07 | 0.28 ±0.0 | 0.27 ±0.0 | 0.28 ±0.0 |
| AIDS | 0.84 ±0.04 | 0.96 ±0.06 | 0.84 ±0.04 | NA | 1.0 ±0.0 | NA | 0.19 ±0.02 | 0.20 ±0.03 | 0.19 ±0.04 |
| ogbg-molhiv | 0.99 +- 0.03 | 0.99 +- 0.04 | 0.99 +- 0.04 | 0.801 ±0.149 | 0.764 ±0.14 | 0.784 ±0.144 | OOM | OOM | OOM |
Stability of counterfactual explainers: Table 8 provides an overview of the stability exhibited among explainer instances trained using three distinct seeds. Notably, we observe a substantial Jaccard index, indicating favorable stability, in the case of RCExplainer and Cf explainers. Conversely, Clear fails to demonstrate comparable stability. These findings align with the outcomes derived from Table 4. Specifically, when RCExplainer and Cf are successful in identifying a counterfactual, the resultant counterfactual graphs are obtained through a small number of perturbations. Consequently, the counterfactual graphs exhibit similarities to the original graph, rendering them akin to one another. However, this trend does not hold true for Clear, as it necessitates a significantly greater number of perturbations.
Similar observations are made concerning stability in the presence of topological noise and various Gnn architectures, owing to the aforementioned reasons. For detailed results, please refer to App. B.
@@ 4.3 Necessity and Reproducibility
Recall the definitions of necessity and sufficiency from § 3. In both cases, we evaluate performance using sufficiency as the metric. Sufficiency, in this context, measures the ratio of graphs for which the Gnn prediction on the residual graph is the same as in the original graph. Following the removal of explanation, we expect sufficiency to be low.
The results are presented in the form of plots in App. C (due to space limitations). While sufficiency is indeed low in necessity, this is not the case in reproducibility. These findings suggest that while current factual explainers effectively explain the model, they do not provide a comprehensive explanation of the underlying data. The fact that the Gnn can regain prediction accuracy when retrained indicates the presence of other signals that the initial factual explanation failed to capture.
@@ 4.4 Feasibility
Counterfactual explanations serve as recourses and are expected to generate graphs that adhere to the feasibility constraints of the pertinent domain. Given that the benchmarked algorithms are agnostic to specific domains, our initial evaluation entails examining whether these algorithms preserve the topological properties of the test set. To achieve this, we compare the number of graphs forming a single connected component in the test set with those in their corresponding counterfactual explanations. Connectedness is a significant aspect of consideration, particularly in domains such as molecules, where disconnected graphs are rare occurrences and might not be meaningful. The results for RCExplainer are presented in Table 9.
| Dataset | Expected Count | Observed Count | -value |
| Mutagenicity | 233.05 | 70 | < 0.00001 |
| Proteins | 11.68 | 12 | 0.93 |
| Mutag | 11.0 | 9.0 | 0.55 |
| AIDS | 17.6 | 8 | 0.02 |
We use RCExplainer for this experiment since it is the most robust and well-performing algorithm among those we have benchmarked. Notably, we observe statistically significant deviations from the expected values in two out of four molecular datasets. This suggests a heightened probability of predicting counterfactuals that do not correspond to feasible molecules. Importantly, this finding underscores a limitation of counterfactual explainers, which has received limited attention within the research community.
@@ 4.5 Visualization-based Analysis
We include visualization based analysis of the explanations in App. D. Due to space limitations, we omit the details here. Our analysis reveal that a statistically good performance do not always align with human judgement indicating an urgent need for datasets annotated with ground truth explanations. Furthermore, the visualization analysis reinforces the need to incorporate feasibility as desirable component in counterfactual reasoning.
5 Concluding Insights
Our benchmarking study has yielded several insights that can streamline the development of explanation algorithms. We summarize the key findings below.
@@ 5 Concluding Insights
Model Explanation vs. Data Explanation: Our experiments on reproducibility (§ 4.3) revealed that even without the factual explanation, the Gnn model predicted accurately on the residual graph. This suggests that explainers only capture specific signals learned by the Gnn and do not encompass all underlying data signals.
@@
We hope that the aforementioned insights would offer new directions for advancing Gnn explainers, allowing researchers to address limitations and enhance the overall quality and interpretability of Gnns.
Limitations. Our study only focuses on perturbation-based instance-specific methods. We hope to replicate the same study on the branches shown in Fig. 1.
@@ A.1 Benchmark Datasets
MUTAG [20] and Mutagenicity [37, 23]: These are graph datasets containing chemical compounds. The nodes represent atoms, and the edges represent chemical bonds. The binary labels depend on the mutagenic effect of the compound on a bacterium, namely mutagenic or non-mutagenic. MUTAG and Mutagenicity datasets contain 188 and 4337 graphs, respectively.
@@
AIDS: [20] This dataset contains small molecules. The nodes and edges are atoms and chemical bonds, respectively. The molecules are classified by whether they are active against the HIV virus or not.
@@
NCI1 [48]: This dataset is derived from cheminformatics and represents chemical compounds as input graphs. Vertices in the graph correspond to atoms, while edges represent bonds between atoms. This dataset focuses on anti-cancer screenings for cell lung cancer, with chemicals labeled as positive or negative. Each vertex is assigned an input label indicating the atom type, encoded using a one-hot-encoding scheme.
@@
IMDB-B [55]: The IMDB-BINARY dataset is a collection of movie collaboration networks, encompassing the ego-networks of 1,000 actors and actresses who have portrayed roles in films listed on IMDB. Each network is represented as a graph, where the nodes correspond to the actors/actresses, and an edge is present between two nodes if they have shared the screen in the same movie. These graphs have been constructed specifically from movies in the Action and Romance genres, which are the class labels.
@@
REDDIT-B [55]: REDDIT-BINARY dataset encompasses graphs representing online discussions on Reddit. Each graph has nodes representing users, connected by edges when either user responds to the other’s comment. The four prominent subreddits within this dataset are IAmA, AskReddit, TrollXChromosomes, and atheism. IAmA and AskReddit are question/answer-based communities, while TrollXChromosomes and atheism are discussion-based communities. Graphs are labeled based on their affiliation with either a question/answer-based or discussion-based community.
@@
GRAPH-SST2 [61]: The Graph-SST2 dataset is a graph-based dataset derived from the SST2 dataset [42], which contains movie review sentences labeled with positive or negative sentiment. Each sentence in the Graph-SST2 dataset is transformed into a graph representation, with words as nodes and edges representing syntactic relationships capturing the sentence’s grammatical structure. The sentiment labels from the original SST2 dataset are preserved, allowing for sentiment analysis tasks using the graph representations of the sentences.
@@
ogbg-molhiv [3]: ogbg-molhiv is a molecule dataset with nodes representing atoms and edges representing chemical bonds. The node features represent various properties of the atoms like chirality, atomic number, formal charge etc. Edge attributes represent the bond type. We study binary classification task on this dataset. The task is to achieve the most accurate predictions of specific molecular properties. These properties are framed as binary labels, indicating attributes like whether a molecule demonstrates inhibition of HIV virus replication or not.
A.2 Factual Explainers: Inductive Methods on Test Set
A.3 Details of Gnn model used for Node Classification
@@
We use the same Gnn model used in CF-GnnExplainer and Cf. Specifically, it is a Graph Convolutional Networks [24] trained on each of the datasets. Each model has 3 graph convolutional layers with 20 hidden dimensions for the benchmark datasets. The non-linearity used is relu for the first two layers and log after the last layer of GCN. The learning rate is 0.01. The train and test data are divided in the ratio 80:20. The accuracy of the Gnn model for each dataset is mentioned in Table J.
@@
A.4 Details of Base Gnn Model for the Graph Classification Task
Our GNN models have an optional parameter for continuous edge weights, which in our case represents explanations. Each model consists of 3 layers with 20 hidden dimensions specifically designed for benchmark datasets. The models provide node embeddings, graph embeddings, and direct outputs from the model (without any softmax function). The output is obtained through a one-layer MLP applied to the graph embedding. We utilize the max pooling operator to calculate the graph embedding. The dropout rate, learning rate, and batch size are set to 0, 0.001, and 128, respectively. The train, validation, and test datasets are divided into an 80:10:10 ratio. The algorithms run for 1000 epochs with early stopping after 200 patience steps on the validation set. The performance analysis of the base GNN models for each graph classification dataset is presented in Table K.
@@ A.4 Details of Base Gnn Model for the Graph Classification Task
A.5 Details of Factual Explainers for the Graph Classification Task
@@
On the other hand, SubgraphX ranks different subgraph explanations based on their scores. We select the top 20 explanations and, for each explanation, compute the subgraph. Then, we enhance the importance of each edge of a particular subgraph by incrementing its score by 1. Finally, we normalize the weights of the edges. This process allows us to obtain continuous explanations as well.
@@ A.6 Codes and Implementation
Table L shows the code bases we have used for the explainers. We have adapted the codes based on our base GNN models. Our repository, https://github.com/Armagaan/gnn-x-bench, includes the adaptations of the methods to our base models.
A.7 Feasibility for Factual Explanations

### v3

@@ 4.2 Stability
We next examine the stability of the explanations against topological noise, model parameters, and the choice of Gnn architecture. In App. C, we present the impact of the above mentioned factors on other metrics of interest such as sufficiency and explanation size. In addition, we also present impact of feature perturbation and topological adversarial attack in App. C.
@@
Insights on factual-stability against topological noise: Fig. 3 illustrates the Jaccard coefficient as a function of the noise volume. Similar to Fig.2, edge selection for the explanation involves a greedy approach that prioritizes the highest score edges. A clear trend that emerges is that inductive methods consistently outperform the transductive methods (such as Cf and GNNExplainer). This is expected since transductive methods lack generalizable capability to unseen data. Furthermore, the stability is worse on denser datasets of IMDB-B since due to the presence of more edges, the search space of explanation is larger. RCExplainer (executed at ) and PGExplainer consistently exhibit higher stability. This consistent performance reinforces the claim that RCExplainer is the preferred factual explainer. The stability of RCExplainer can be attributed to its strategy of selecting a subset of edges that is resistant to changes, such that the removal of these edges significantly impacts the prediction made by the remaining graph Bajaj et al. (2021). PGExplainer also incorporates a form of inherent stability within its framework. It builds upon the concept introduced in GNNExplainer through the assumption that the explanatory graph can be modeled as a random Gilbert graph, where the probability distribution of edges is conditionally independent and can be parameterized. This generic assumption holds the potential to enhance the stability of the method. Conversely, TagExplainer exhibits the lower stability than RCExplainer and PGExplainer, likely due to its reliance solely on gradients in a task-agnostic manner Xie et al. (2022). The exclusive reliance on gradients makes it more susceptible to overfitting, resulting in reduced stability.
@@
Insights on factual-stability against explainer instances: Table 7 presents the stability of explanations provided across three different explainer instances on the same black-box Gnn. A similar trend is observed, with RCExplainer remaining the most robust method, while GnnExplainer exhibits the least stability. For GnnExplainer, the Jaccard coefficient hovers around , indicating significant variance in explaining the same Gnn. Although the explanations change, their quality remains stable (as evident from small standard deviation in Fig. 2). This result indicates that multiple explanations of similar quality exist and hence a single explanation fails to complete explain the data signals. This component is further emphasized when we delve into reproducibility (§ 4.3).
@@
Insights on factual-stability against Gnn architectures: Finally, we explore the stability of explainers across different Gnn architectures in Table 8, which has not yet been investigated in the existing literature. For each combination of architectures, we assess the stability by computing the Jaccard coefficient between the explained predictions of the indicated Gnn architecture and the default Gcn model. One notable finding is that the stability of explainers exhibits a strong correlation with the dataset used. Specifically, in five out of six datasets, the best performing explainer across all architectures is unique. However, it is important to highlight that the Jaccard coefficients across architectures consistently remain low indicating stability against different architectures is the hardest objective due to the variations in their message aggregating schemes.
@@ 4.3 Necessity and Reproducibility
We aim to understand the quality of explanations in terms of necessity and reproducibility. The results are presented in App. D and E. Our findings suggest that necessity is low but increases with the removal of more explanations, while reproducibility experiments reveal that explanations do not provide a comprehensive explanation of the underlying data, and even removing them and retraining the model can produce a similar performance to the original Gnn model.
@@ 4.4 Feasibility
Counterfactual explanations serve as recourses and are expected to generate graphs that adhere to the feasibility constraints of the pertinent domain. We conduct the analysis of feasibility on molecular graphs. It is rare for molecules to be constituted of multiple connected components Vismara & Laurenço (2000). Hence, we study the distribution of molecules that are connected in the original dataset and its comparison to the distribution in counterfactual recourses. We measure the -value of this deviation. App. A.7 presents the results.
@@ 4.5 Visualization-based Analysis
We include visualizations of the explanations in App. F. Our analysis reveals that a statistically good performance does not always align with human judgment indicating an urgent need for datasets annotated with ground truth explanations. Furthermore, the visualization analysis reinforces the need to incorporate feasibility as a desirable component in counterfactual reasoning.
5 Concluding Insights and Potential Solutions
Our benchmarking study has yielded several insights that can streamline the development of explanation algorithms. We summarize the key findings below (please also see the App. K for our recommendations of explainer for various scenarios).
@@ 5 Concluding Insights and Potential Solutions
Model Explanation vs. Data Explanation: Reproducibility experiments (§ 4.3) revealed that retraining with only factual explanations cannot reproduce the predictions fully. Furthermore, even without the factual explanation, the Gnn model predicted accurately on the residual graph. This suggests that explainers only capture specific signals learned by the Gnn and do not encompass all underlying data signals.
@@
Potential Solutions: The aforementioned insights raise important shortcomings that require further investigation. Below, we explore potential avenues of research that could address these limitations.
Feasible recourses through counterfactual reasoning: Current counterfactual explainers predominantly concentrate on identifying the shortest edit path that nudges the graph toward the decision boundary. This design inherently neglects the feasibility of the proposed edits. Therefore, it is imperative to explicitly address feasibility as an objective in the optimization function. One potential solution lies in the vibrant research field of generative modeling for graphs, which has yielded impressive results Goyal et al. (2020); You et al. (2018); Vignac et al. (2023). Generative models, when presented with an input graph, can predict its likelihood of occurrence within a domain defined by a set of training graphs. Integrating generative modeling into counterfactual reasoning by incorporating likelihood of occurrence as an additional objective in the loss function presents a potential remedy.
Ante-hoc explanations for stability and reproducibility: We have observed that if the explanations are removed and the Gnn is retrained on the residual graphs, the Gnn is often able to recover the correct predictions from our reproducibilty experiments. Furthermore, the explanation exhibit significant instability in the face of minor noise injection. This incompleteness of explainers and instability is likely a manifestation of their post-hoc learning framework, wherein the explanations are generated post the completion of Gnn training. In this pipeline, the explainers have no visibility to how the Gnn would behave to perturbations on the input data, initialization seeds, etc. Potential solutions may lie on moving to an ante-hoc paradigm where the Gnn and the explainer are jointly trained Kosan et al. (2023); Miao et al. (2022); Fang et al. (2023).
These insights, we believe, open new avenues for advancing Gnn explainers, empowering researchers to overcome limitations and elevate the overall quality and interpretability of Gnns.
6 Acknowledgements
Samidha Verma acknowledges the generous grant received from Microsoft Research India to sponsor her travel to ICLR 2024. Additionally, this project was partially supported by funding from the National Science Foundation under grant #IIS-2229876.
@@ A.1 Benchmark Datasets
BA-SHAPES Ying et al. (2019b): The base graph in this dataset is a Barabasi-Albert (BA) graph. The dataset includes house-shaped structures composed of 5 nodes (Figure D (c)). Non-motif nodes are assigned class , while nodes at the top, middle, and bottom of the motif are assigned classes , , and , respectively.
@@
MUTAG Ivanov et al. (2019) and Mutagenicity Riesen & Bunke (2008); Kazius et al. (2005): These are graph datasets containing chemical compounds. The nodes represent atoms, and the edges represent chemical bonds. The binary labels depend on the mutagenic effect of the compound on a bacterium, namely mutagenic or non-mutagenic. MUTAG and Mutagenicity datasets contain 188 and 4337 graphs, respectively.
@@
AIDS: Ivanov et al. (2019) This dataset contains small molecules. The nodes and edges are atoms and chemical bonds, respectively. The molecules are classified by whether they are active against the HIV virus or not.
@@
NCI1 Wale et al. (2008): This dataset is derived from cheminformatics and represents chemical compounds as input graphs. Vertices in the graph correspond to atoms, while edges represent bonds between atoms. This dataset focuses on anti-cancer screenings for cell lung cancer, with chemicals labeled as positive or negative. Each vertex is assigned an input label indicating the atom type, encoded using a one-hot-encoding scheme.
@@
IMDB-B Yanardag & Vishwanathan (2015): The IMDB-BINARY dataset is a collection of movie collaboration networks, encompassing the ego-networks of 1,000 actors and actresses who have portrayed roles in films listed on IMDB. Each network is represented as a graph, where the nodes correspond to the actors/actresses, and an edge is present between two nodes if they have shared the screen in the same movie. These graphs have been constructed specifically from movies in the Action and Romance genres, which are the class labels.
@@
REDDIT-B Yanardag & Vishwanathan (2015): REDDIT-BINARY dataset encompasses graphs representing online discussions on Reddit. Each graph has nodes representing users, connected by edges when either user responds to the other’s comment. The four prominent subreddits within this dataset are IAmA, AskReddit, TrollXChromosomes, and atheism. IAmA and AskReddit are question/answer-based communities, while TrollXChromosomes and atheism are discussion-based communities. Graphs are labeled based on their affiliation with either a question/answer-based or discussion-based community.
@@
GRAPH-SST2 Yuan et al. (2022): The Graph-SST2 dataset is a graph-based dataset derived from the SST2 dataset Socher et al. (2013), which contains movie review sentences labeled with positive or negative sentiment. Each sentence in the Graph-SST2 dataset is transformed into a graph representation, with words as nodes and edges representing syntactic relationships capturing the sentence’s grammatical structure. The sentiment labels from the original SST2 dataset are preserved, allowing for sentiment analysis tasks using the graph representations of the sentences.
@@
ogbg-molhiv Allamanis et al. (2018): ogbg-molhiv is a molecule dataset with nodes representing atoms and edges representing chemical bonds. The node features represent various properties of the atoms like chirality, atomic number, formal charge etc. Edge attributes represent the bond type. We study binary classification task on this dataset. The task is to achieve the most accurate predictions of specific molecular properties. These properties are framed as binary labels, indicating attributes like whether a molecule demonstrates inhibition of HIV virus replication or not.
A.2 Details of Gnn model used for Node Classification
@@
We use the same Gnn model used in CF-GnnExplainer and Cf. Specifically, it is a Graph Convolutional Networks Kipf & Welling (2016) trained on each of the datasets. Each model has 3 graph convolutional layers with 20 hidden dimensions for the benchmark datasets. The non-linearity used is relu for the first two layers and log after the last layer of GCN. The learning rate is 0.01. The train and test data are divided in the ratio 80:20. The accuracy of the Gnn model for each dataset is mentioned in Table I.
@@
A.3 Details of Base Gnn Model for the Graph Classification Task
Our GNN models have an optional parameter for continuous edge weights, which in our case represents explanations. Each model consists of 3 layers with 20 hidden dimensions specifically designed for benchmark datasets. The models provide node embeddings, graph embeddings, and direct outputs from the model (without any softmax function). The output is obtained through a one-layer MLP applied to the graph embedding. We utilize the max pooling operator to calculate the graph embedding. The dropout rate, learning rate, and batch size are set to 0, 0.001, and 128, respectively. The train, validation, and test datasets are divided into an 80:10:10 ratio. The algorithms run for 1000 epochs with early stopping after 200 patience steps on the validation set. The performance analysis of the base GNN models ( Kipf & Welling (2016); Veličković et al. (2018); Xu et al. (2019); Hamilton et al. (2017)) for each graph classification dataset is presented in Table J.
@@ A.3 Details of Base Gnn Model for the Graph Classification Task
A.4 Details of Factual Explainers for the Graph Classification Task
@@
On the other hand, SubgraphX ranks different subgraph explanations based on their scores. We select the top 20 explanations and, for each explanation, compute the subgraph. Then, we enhance the importance of each edge of a particular subgraph by incrementing its score by 1. Finally, we normalize the weights of the edges. This process allows us to obtain continuous explanations as well. Moreover, since SubgraphX employs tree search, its scalability is limited when dealing with large graphs. For instance, in the Mutagenicity dataset, obtaining explanations for 435 graphs requires approximately 26.5 hours. To address this challenge, we restricted our analysis to test graphs when calculating explanations using SubgraphX. It is important to include this disclaimer, working on only subset of graphs may introduce potential biases or noises in the results.
A.5 Factual Explainers: Inductive Methods on Test Set
@@ A.6 Codes and Implementation
Table K shows the code bases we have used for the explainers. We have adapted the codes based on our base GNN models. Our repository, https://github.com/Armagaan/gnn-x-bench/, includes the adaptations of the methods to our base models.
| Method | Repository |
| PGExplainer Luo et al. (2020) | https://github.com/LarsHoldijk/RE-ParameterizedExplainerForGraphNeuralNetworks/ |
| TAGExplainer Xie et al. (2022) | https://github.com/divelab/DIG/tree/main/dig/xgraph/TAGE/ |
| CF Tan et al. (2022) | https://github.com/chrisjtan/gnn_cff |
| RCExplainer Bajaj et al. (2021) | https://developer.huaweicloud.com/develop/aigallery/notebook/detail?id=e41f63d3-e346-4891-bf6a-40e64b4a3278 |
| GNNExplainer Ying et al. (2019b) | https://github.com/LarsHoldijk/RE-ParameterizedExplainerForGraphNeuralNetworks/ |
| GEM Lin et al. (2021b) | https://github.com/wanyu-lin/ICML2021-Gem/ |
| SubgraphX Yuan et al. (2021) | https://github.com/divelab/DIG/tree/main/dig/xgraph/SubgraphX |
| Dataset | RCExplainer | Cf | ||||
| Expected Count | Observed Count | -value | Expected Count | Observed Count | -value | |
| Mutagenicity | 233.05 | 70 | < 0.00001 | 206.65 | 0 | < 0.00001 |
| Mutag | 11 | 9 | 0.55 | 4 | 1 | 0.13 |
| AIDS | 17.6 | 8 | < 0.00001 | 1.76 | 0 | 0.0001 |
A.7 Feasibility
Counterfactual explanations: As shown in Table L, we observe statistically significant deviations from the expected values in two out of three molecular datasets. This suggests a heightened probability of predicting counterfactuals that do not correspond to feasible molecules. This finding underscores a limitation of counterfactual explainers, which has received limited attention within the research community.